please help

An HCl solution has a pH = 4. If you dilute 10 mL of the solution to 100mL, the final pH will be:
Answer #1

SInce we know that

[H+] for 10 ml = 10-4 = 0.0001 = M1

now we know that from dilution law

M1V1 = M2V2

or M2 = 0.0001*10 / 100= 0.00001

now

pH of diluted solution = -log[0.00001] = 5

the final pH will be 5

Please give me thumbs up